# On the Scalability of GNNs for Molecular Graphs

### v3

@@ 4.4 Finetuning and Probing
Finetuning. Since our training setup consists of multiple tasks and our architectures incorporate multiple task heads, we need to identify a finetuning module after which the remaining pretraining architecture is removed and replaced by a newly initialized MLP: the finetuning head. As all downstream tasks discussed above reside on the graph level, our main choice is the graph output network, the MLP that processes features after being aggregated from the node to the graph level, and further feeds into the task heads for graph-level tasks. Intuitively, this layer’s output representations have benefited most from pretraining on diverse data and tasks, as it feeds directly into the various task heads. We further investigate the effect of choosing layers of the tasks heads as finetuning module to potentially leverage synergies between specific pretraining and downstream tasks. As all downstream tasks are on the graph level, we trim the architecture by removing parts related to node level tasks and unused task heads.
Probing and Fingerprinting. Similar to finetuning, we employ probing using an MLP as a suitable strategy for obtaining general representations on downstream tasks. For probing, the base model is kept frozen and only the new layers are trained. This allows the training procedure to focus the gradient on newly added parameters, resulting in task-specific head layers. In the case of large model sizes, running features through the frozen base model is expensive with respect to memory and compute. We tackle this bottleneck by caching hidden representations on disk and reusing them during probing. Since the gradient does not impact parameters of the base model, fingerprints remain unchanged yielding an optimized strategy for downstream tasks capable of parallelization across multiple inexpensive devices. In this work, similar to the finetuning setup, we extract fingerprints from the task head MLPs of graph level tasks, and from the graph output network, the MLP that directly feeds in to the task heads.
@@ 5 Experiments
Figure 2: Effect of scaling different design choices (columns) to model evaluation performance (rows). The standardized mean is calculated as mean of standardized scores for every task in a dataset group, i.e., a mean and standard deviation per task were calculated based on all our models in this study (signs of tasks with lower is better metrics were flipped).
@@ 5.1 Scaling Trends for Pretraining
MPNN++ vs Transformer. MPNN++ models are more parameter efficient as they perform better with small width and depth compared to Transformers. They are also data efficient as they perform significantly better for the quantum PCQM4M_* tasks when sub-sampling the datasets, although smaller differences are observed for the remaining (biological) datasets. Transformers being “data-hungry” is consistent with recent works in other domains such as natural language and computer vision (Radford et al., 2018; Alayrac et al., 2022; Galkin et al., 2023). The hybrid GPS++ architecture seems to benefit from the MPNN++ expressivity in low-parameter regimes, while also exhibiting a similar molecular scaling to the Transformer in low-data regimes. Finally, we notice that MPNN++ models are more affected by depth, an unsurprising outcome considering that, contrarily to Transformers, their receptive field depend on the number of layers.
@@
Depth scaling. Similar to width, depth of GNN models plays an important role in the dataset fit during test time. Deeper models with larger layers capture intricate aspects of the data resulting in 12.5% improvement in test error. However, performance plateaus at around 8-16 layers for quantum datasets, while could be mitigated by larger datasets. For PCBA_1328, the performance continues to increase.
@@
Detailed scaling law analysis. We provide more detailed analysis of the observed scaling trends in terms of Equations 1 and 2 in Appendix F, observing scaling laws similar to those in other domains (Kaplan et al., 2020).
@@ 5.2 Scaling Trends on Downstream Tasks
Figure 3: Finetuning and probing performance of pretrained MPNN++ models across various widths on the Polaris benchmark. Darker green shades denote higher/desirable metric values. Larger models tend to perform better across different metrics on unseen tasks. Spearman correlation values closer to 1 indicate that predictive performance correlates with larger model sizes.
@@
We now evaluate scaling of models when finetuning and probing on downstream tasks. As detailed in Section 4.4, all weights are tuned in the case of finetuning, while the pretrained model is frozen when fingerprinting followed by probing. Due to the large number of downstream tasks spread across 38 tasks, we limit our evaluation to probing for most experiments, except for MPNN++ where we also finetune the model.
@@
To summarize scaling trends, we compute the Spearman’s rank correlation coefficient (Schober et al., 2018) between model performance on a given metric and the scale of model/data used. The correlation is given by a value in the range $[-1,1]$ , with a value of $1$ indicating perfect scaling (i.e., a larger model or dataset yields better downstream task performance), $-1$ indicating imperfect scaling (i.e., a smaller model or dataset would be preferred) and 00 indicating no correlation. We note that this evaluation scheme, although statistical, aims to answer the question What kind of design decisions are necessary to build a foundational model for molecular representations?
@@
MPNN++ vs Transformer. For probing on downstream tasks, we study the effect of architecture choices of width, depth, and number of molecules. We find that Transformers benefit from increased width on downstream tasks compared to GPS++ and MPNN++. Surprisingly, despite the number of molecules having a stronger impact on all model’s performance, it only slightly impacts the downstream performance of all models, with a small benefit for MPNN++. Finally, Transformer is the only model with a small positive trend for depth scaling, while GPS++ and MPNN++ show close to no trend.
@@
Figure 4: Comparison of our MolGPS foundation model (that combines fingerprints from the MPNN++, Transformer and hybrid GPS++ model) to the SOTA across TDC, Polaris, and MoleculeNet benchmarks. For TDC, a group of 8 different models establishes SOTA across the 22 tasks. MolGPS establishes new SOTA on $12/22$ TDC tasks and on all but two tasks among Polaris and MoleculeNet.
@@
Width scaling. We evaluate width scaling on Polaris and TDC datasets in Figures 3 and 9 (Appendix E.1). We observe linear scaling trends for MPNN++ on all Polaris datasets, with an average Spearman correlation of 0.91 for probing and 0.85 for finetuning. On TDC, a similar trends are observed (on average 0.69 for probing and 0.72 for finetuning) with a strong correlation of $>\!0.75$ for $15/22$ datasets during probing and $17/22$ during finetuning. These results strongly indicate the benefits of larger pretrained GNNs for downstream tasks, a result consistent with prior findings in scaling of large models (Kaplan et al., 2020). Similarly, Transformer and GPS++ show strong positive scaling trends with model width.
@@
Depth scaling. We evaluate the scaling of depth of MPNN++, GPS++ and Transformer models on the Polaris and TDC benchmarks in Figures 10 and 11 (Appendix E.2). For probing on Polaris, we observe weak positive trends, with average scaling Spearman correlations of $0.47$, $0.55$, and $0.50$, respectively. We see weaker average correlations on TDC, being slightly negative for MPNNN++ and GPS++ and best for Transformer with $0.27$. However, finetuning MPNN++ achieves a respectable correlation of $0.33$. While some datasets strongly benefit from deeper networks, others strongly deteriorate with no clear pattern observable for the TDC datasets. We conjecture that degradation with depth is related to the oversmoothing issue described in Appendix B. Certain molecular properties can be well predicted only from small local substructures, hence eliminating the need for long-range interactions that deeper networks enable.
@@
We further conducted a study to determine the importance of the pretraining data in two ways. Firstly, we repeat pretraining of the models without specific pretraining datasets (dataset ablation). Secondly, we probe models specifically from certain task head MLPs compared to the base GNN models (task-head ablations).
Dataset ablation. Observing the dataset ablations in Figure 16 and 17 (Appendix E.5), we see that PCBA_1328 is the most important pretraining dataset for downstream task performance while L1000_* actually hurts the performance on certain tasks. It will therefore prove beneficial to pretrain without the L1000_* datasets as we will see later.
Task-head ablations. We further tested the effect of probing from different layers of the task heads rather than the graph output network. Results are shown in Figure 18 (Appendix E.6). While overall, the graph output network leads to best performance and correlation, the representation after the first layer of the PCBA_1328 task head performs strikingly well for some tasks, possibly due to synergies from pretraining on a bio assay dataset. This suggests probing approaches using combinations of fingerprints could further improve results. On the other hand, the layers from the PCQM4M_G25 dataset perform poorly, which is intuitive as this pretraining task is dissimilar to the downstream task.
@@ 5.3 Towards a final foundation model
We now explain how the above findings can be pieced together to develop a highly competitive graph foundation model. Apart from scaling the model width, we found two other design choices that had a major impact on the performance for the various downstream tasks. We additionally report results on the MoleculeNet benchmark (Wu et al., 2018) here in addition to the previously used TDC and Polaris benchmarks.
@@
Multi-fingerprint probing. Our previous task-head ablation study suggested that different fingerprints may be optimal for probing depending on the downstream task. As a result we further experimented with probing on combinations of multiple fingerprints extracted at different layers of a pretrained model, which improves performance on downstream tasks. Results can be further enhanced by combining fingerprints from multiple pretrained models.
@@
Pretraining without L1000. Additionally, based on our observation in the dataset ablation, we pretrained new models without the L1000_* pretraining tasks, which leads to performance improvements across all scales. We hypothesize this is due to the challenging signal-to-noise ratio for those particular tasks, as also pointed out in the literature (Tong et al., 2023).
[…]
@@
Figure 5: Comparison of our GPS++ probing (that leverages multiple fingerprints) and our MolGPS foundation model (that leverages fingerprints from MPNN++, Transformer, and the hybrid GPS++) to various baselines across TDC benchmark collection using an aggregated metric (average standardized mean of the leaderboard).
@@
MolGPS. We now introduce a final graph foundation model that leverages the various findings of this paper. MolGPS inherits many architecture design choices from the General, Powerful, Scalable Graph Transformer method (Rampášek et al., 2022) and can be used to navigate the molecular space. MolGPS combines fingerprints from 1B MPNN++, Transformer, and GPS++ models that have been pretrained without the L1000 dataset, followed by an MLP. Figure 4 compares this model across the TDC, Polaris and MoleculeNet benchmark collections to the current SOTA for each task and to MolE. MolGPS yields by far the strongest downstream task results, outperforming MolE in $19/22$ TDC tasks and establishing SOTA performance on $12/22$ TDC tasks. This makes MolGPS the model with the most SOTA entries in the TDC leaderboard followed by MapLight+GNN (Notwell & Wood, 2023) that established SOTA on 5 TDC tasks and 7 other methods that are SOTA for at least one TDC task. [^3] We report scores from the TDC leaderboards extracted on March 15, 2024. SOTA on TDC is established by a group of 8 different models, namely Chemprop-RDKit (Yang et al., 2019), MapLight, MapLight+GNN (Notwell & Wood, 2023), BaseBoosting (Huang12 et al., ), CFA (Quazi et al., 2023), SimGCN, ZairaChem (Oselusi et al., 2024) and ContextPred (Zhou et al., 2020). Similarly, when compared to previous best methods on the Polaris and MoleculeNet benchmarks, we observe that our model is significantly better (often by large margins) for all but two downstream tasks. We primarily attribute the large-scale success of our model to width scaling up to 1B parameters, suggesting us to scale even further.
In Figure 5, we also report the aggregated score on the TDC benchmark for MolGPS, which almost reaches the aggregated SOTA performance, which is remarkable recalling again that this score is a derived from the best scoring method per task of the benchmark collection, while we use a single method for all tasks.
@@ 6 Conclusion
In this paper, we studied the scalability of GNN models including message-passing networks, graph Transformers and hybrid architectures on the largest public collection of 2D molecules for the tasks of molecular property prediction. We showed major performance gains from the growing amount of parameters, data and compute, both on the original test set and on downstream finetuning. Importantly, our models benefit tremendously from the increasing scale of width, number of molecules, and number of labels. Our largest 1B parameter models, including MPNN++, Transformer, and GPS++, continue to scale favourably. More importantly, we demonstrate a consistent performance improvement on downstream property prediction tasks via finetuning and probing. Finally, we derive MolGPS, a powerful foundational model based on a multi-fingerprint probing approach that can be used to navigate the chemical space, establishing state-of-the-art on 26 out of 38 highly competitive downstream tasks. We hope that our work paves the way for the development of foundational GNNs and new architectures with applications in pharmaceutical advancements and drug discovery.
6.1 Future Work
@@
While our study demonstrates the benefits of increasing number of parameters far greater than prior work, there are still orders of magnitude before we reach a general-purpose foundational model of molecules. Our analysis is restricted to the effect of number of parameters and molecules during pretraining and finetuning stages. Future work would aim to uncover additional aspects of GNN training such as the increasing complexity of aggregation functions and their effect on scaling properties.
7 Impact Statement
@@
This paper presents work where we aim to build accurate foundational models for molecular property prediction. Similar to other areas and applications of machine learning, our foundational models can have significant societal impact, both positive and negative. We believe the analysis presented in the paper will be most useful in designing new compounds with different molecular geometries and properties. We thus foresee positive impacts of GNNs in areas of drug discovery, pharmaceutical advancements and tackling rare diseases by studying their molecular configurations. The scaling of different GNN model choices can be potentially useful in simultaneously predicting and optimizing novel molecular compositions. Models could also be used to assess unknown properties of drugs and help prevent undesired side-effects. On the other hand, such models could also be used for harmful purposes such as developing chemical weapons and biohazards. We note that the usage of GNN models for such applications is less likely.
Appendix A Experimental Details
A.1 Pretraining

### v4

@@ 3.4 Finetuning and Probing
Finetuning. Since our training setup consists of multiple tasks and our architectures incorporate multiple task heads, we need to identify a finetuning module after which the remaining pretraining architecture is removed and replaced by a newly initialized MLP: the finetuning head. As all downstream tasks discussed above reside on the graph level, our main choice is the graph output network, the MLP that processes features after being aggregated from the node to the graph level, and further feeds into the task heads for graph-level tasks. Intuitively, this layer’s output representations have benefited the most from pretraining on diverse data and tasks, as it feeds directly into the various task heads. We further investigate the effect of choosing layers of the tasks heads as finetuning module to potentially leverage synergies between specific pretraining and downstream tasks. As all downstream tasks are on the graph level, we trim the architecture by removing parts related to node level tasks and unused task heads.
Probing and Fingerprinting. Similar to finetuning, we employ probing using an MLP as a strategy for solving new downstream tasks. For probing, the base model is kept frozen and only the new layers are trained. This allows the training procedure to focus the gradient on newly added parameters, resulting in task-specific probing head layers. In the case of large model sizes, extracting embeddings (so-called fingerprints) from the frozen base model is expensive with respect to memory consumption and compute. We tackle this bottleneck by caching fingerprints on disk and reusing them during probing. Since the gradient does not impact parameters of the base model, fingerprints remain unchanged yielding an optimized strategy for downstream tasks capable of parallelization across multiple inexpensive devices. In this work, similar to the finetuning setup, we extract fingerprints from the task head MLPs of graph level tasks, and from the graph output network, the MLP that directly feeds in to the task heads.
@@ 4.1 Scaling Trends for Pretraining
Figure 2: Effect of scaling different scaling types (columns) to test performance (rows). The standardized mean is calculated as mean of standardized scores for every task in a dataset group, i.e., a mean and standard deviation per task were calculated based on all our models in this study (signs of tasks with lower is better metrics were flipped).
@@
MPNN++ vs Transformer. MPNN++ models are more parameter efficient as they perform better with small width and depth compared to Transformers. They are also data efficient as they perform significantly better for the quantum PCQM4M_* tasks when sub-sampling the datasets, although smaller differences are observed for the remaining (biological) datasets. Transformers being “data-hungry” is consistent with recent works in domains such as natural language and computer vision [50, 2, 17]. The hybrid GPS++ seems to benefit from the MPNN++ expressivity in low-parameter regimes, while also exhibiting a similar molecular scaling to the Transformer in low-data regimes. Finally, we notice that MPNN++ models are more affected by depth, an unsurprising outcome considering that, contrarily to Transformers, their receptive field depend on the number of layers.
@@
Depth scaling. Similar to width, depth of GNN models plays an important role in the dataset fit in test time. Deeper models with larger layers capture intricate aspects of the data resulting in 12.5% improvement in test error. However, performance plateaus at around 8-16 layers for quantum datasets. For PCBA_1328, the performance continues to increase.
@@
Detailed scaling law analysis. We provide detailed analysis of the observed scaling trends in terms of Equations 1 and 2 in Appendix G, observing scaling laws similar to those in other domains [28].
@@ 4.2 Scaling Trends on Downstream Tasks
We now evaluate scaling of models when finetuning and probing on downstream tasks. As detailed in Section 3.4, all weights are tuned in the case of finetuning, while the pretrained model is frozen when fingerprinting followed by probing.
Figure 3: Finetuning and probing performance of pretrained MPNN++ models of different width on the Polaris benchmark. Darker green shades denote better metric values. Larger models tend to perform better on unseen tasks. Spearman correlation values closer to 1 indicate that predictive performance correlates with larger model sizes.
Due to the large number of downstream tasks spread across 38 tasks, we limit our evaluation to probing for most experiments, except for MPNN++ where we also finetune the model.
@@
To summarize scaling trends, we compute the Spearman’s rank correlation coefficient [58] between model performance on a given metric and the scale of model/data used. The correlation is given by a value in the range $[-1,1]$ , with a value of $1$ indicating perfect scaling (i.e., a larger model or dataset yields better downstream task performance), $-1$ indicating imperfect scaling (i.e., a smaller model or dataset would be preferred) and 00 indicating no correlation. We note that this evaluation scheme, although statistical, aims to answer an important question: What kind of design decisions are necessary to build a foundational model for molecular representations?
@@
MPNN++ vs. Transformer. For probing on downstream tasks, we study the effect of architecture choices of width, depth, and number of molecules. We find that Transformers benefit more from increased width on downstream tasks compared to GPS++ and MPNN++ as seen in Figure 2 (bottom two columns). Despite the number of molecules having a stronger impact on all model’s performance, it only slightly impacts the downstream performance of all models, with a small benefit for MPNN++. Finally, Transformer is the only model with a small positive trend for depth scaling, while GPS++ and MPNN++ show close to no trend (Figure 2).
@@
Figure 4: Comparison of our MolGPS foundation model (that combines fingerprints from the MPNN++, Transformer and hybrid GPS++ model) to the SOTA across TDC, Polaris, and MoleculeNet benchmarks. SOTA refers to the maximum value for each dataset. MolGPS establishes new SOTA on $11/22$ TDC tasks and on all but one task among Polaris and MoleculeNet.
@@
Width scaling. We evaluate width scaling on Polaris and TDC datasets in Figure 3 (and Figure 9 provided in the appendix). We observe linear scaling trends for MPNN++ on all Polaris datasets, with an average Spearman correlation of 0.91 for probing and 0.85 for finetuning. On TDC, a similar trends are observed (on average 0.69 for probing and 0.72 for finetuning) with a strong correlation of $>\!0.75$ for $15/22$ datasets during probing and $17/22$ during finetuning. These results strongly indicate the benefits of larger pretrained GNNs for downstream tasks, a result consistent with prior findings in scaling of large models [28]. Similarly, Transformer and GPS++ show strong positive scaling trends with model width.
@@
Depth scaling. We evaluate the scaling of depth of MPNN++, GPS++ and Transformer models on the Polaris and TDC benchmarks in Figures 10 and 11 (Appendix E.2). For probing on Polaris, we observe weak positive trends, with average scaling Spearman correlations of $0.47$, $0.55$, and $0.50$, respectively. We see weaker average correlations on TDC, being slightly negative for MPNNN++ and GPS++ and best for Transformer with $0.27$. However, finetuning MPNN++ achieves a respectable correlation of $0.33$. While some datasets strongly benefit from deeper networks, others strongly deteriorate with no clear pattern observable for the TDC datasets. We conjecture that degradation with depth is related to the oversmoothing issue described in Appendix C. Certain molecular properties can be well predicted only from small local substructures, hence eliminating the need for long-range interactions that deeper networks enable.
@@
Dataset ablation. We further conducted a study to determine the importance of the pretraining data in two ways. Firstly, we repeat pretraining of the models without specific pretraining datasets (dataset ablation). Secondly, we probe models specifically from certain task head MLPs compared to the base GNN models (task-head ablations).
Observing the dataset ablations in Figure 16 and 17 (Appendix E.5), we see that PCBA_1328 is the most important pretraining dataset for downstream task performance while L1000_* actually hurts the performance on certain tasks. It will therefore prove beneficial to pretrain without the L1000_* datasets as we will see later.
Task-head ablations. We further tested the effect of probing from different layers of the task heads rather than the graph output network. Results are shown in Figure 18 (Appendix E.6). While overall, the graph output network leads to best performance and correlation, the representation after the first layer of the PCBA_1328 task head performs strikingly well for some tasks, possibly due to synergies from pretraining on bioassays. This suggests probing approaches using combinations of fingerprints could further improve results. On the other hand, the layers from the PCQM4M_G25 dataset perform poorly, which is intuitive as this pretraining task is dissimilar to the downstream task.
Probing vs. finetuning. So far, we have considered finetuning and probing side-by-side, establishing both as effective strategies for tackling downstream tasks. However, given the relatively similar performance and the significantly higher computational cost of finetuning, we find probing to be the overall more advantageous approach. Another major benefit of probing is the ability to leverage multi-level information from the pretrained GNN as investigated in our task head ablation study above. We recall that our pretraining is based on a supervised multi-task learning approach. As a result, different task heads capture task-specific information, while earlier layers that feed into the task heads carry more general information. When we combine fingerprints from various layers, we can think of taking into account knowledge from several “experts”.
@@ 4.3 Towards a Final Foundation Model
We now explain how the above findings can be pieced together to develop MolGPS, a powerful graph foundation model. Apart from scaling the model width, we found two other design choices with a major impact on the performance for the various downstream tasks. We report results on the MoleculeNet benchmark [71] here in addition to the previously used TDC and Polaris benchmarks.
@@
Multi-fingerprint probing. Our previous task-head ablation study suggested that different fingerprints may be optimal for probing depending on the downstream task. As a result, we choose probing (instead of finetuning) and further experiment with combinations of multiple fingerprints extracted at different layers of a pretrained model, which improves performance on downstream tasks. Moreover, performance can be further enhanced by combining fingerprints from multiple pretrained models.
@@
Pretraining without L1000. Additionally, based on our observation in the dataset ablation, we pretrained new models without the L1000_* pretraining tasks, which leads to performance improvements across all scales. We hypothesize this is due to the challenging signal-to-noise ratio for those particular tasks, as also pointed out in the literature [64].
@@
Figure 5: Comparison of our MPNN++ probing (that leverages multiple fingerprints; with and without additional phenomics pretraining) and MolGPS (that leverages fingerprints from MPNN++, Transformer and GPS++) to various baselines across TDC benchmark collection using an aggregated metric.
In Figure 5, we present multi-fingerprint probing results of the MPNN++ model. We report an aggregated metric across the TDC benchmark collection (normalized performance222 The normalized performance of a model on a benchmark collection is calculated by first deriving the z-scores of the model’s performance per task relative to the leaderboard. We flip the sign of z-scores for tasks that are “the lower the better“ and then average the z-score across the tasks.). We observe a strictly positive scaling trend up to 1B parameters, clearly outperforming TDC Baseline (the normalized score across the best models per task reported in [26]) and the MolE foundational model [43], a gold standard for molecular property prediction (including a model variant with only self-supervised pretraining). The 1B MPNN++ only slightly trails MapLight+GNN [46], the best-performing single model on the TDC benchmark. Surprisingly, we were unable to further scale the model to the 3B parameter regime, likely due to limitations of our pretraining data mix.
Integrating phenomics into pretraining. To improve our pretraining data mix, we experimented with an additional data type for model pretraining that comes from phenomic imaging. The dataset contains  6k labels for  500k molecules that were derived from phenomic imaging [8] of cells perturbed with either a dose of a compound or a gene knockout. We conducted a similarity analysis between the obtained images (represented by vector embeddings, e.g., similar to He et al. [24]) subject to a compound perturbation on one side, and images subject to a gene perturbation on the other side. The pretraining task is to predict for each compound if it has a phonemically visible similarity to a gene knockout (indicating a biological relationship).
Adding phenomics data to our pretraining data mix (i.e., PCBA_1328 and PCQM4M), improved our downstream task performance across the board. Comparing the scaling trends in Figure 5, MPNN++ with phenomics pretraining exhibits a significant vertical upwards shift compared to the original MPNN++. Notably, we were also able to extend our scaling study to the 3B parameter regime. While we were previously unable to extend the scaling trend, MPNN++ with phenomics maintains a positive scaling trend.
@@
MolGPS. We introduce a final graph foundation model that leverages the various findings of this paper. MolGPS inherits many architecture design choices from the General, Powerful, Scalable Graph Transformer method [53] and can be used to navigate the molecular space. Our 1B and 3B MolGPS combines fingerprints from MPNN++, Transformer, and GPS++ models (of scale 1B and 3B, respectively) that have been pretrained with the new phenomics data and without the L1000 dataset, followed by a specialized probing MLP. Figure 4 compares this model across the TDC, Polaris and MoleculeNet benchmark collections to the current SOTA for each task and to MolE. MolGPS yields by far the strongest downstream task results, outperforming MolE in $21/22$ TDC tasks and establishing SOTA performance on $11/22$ TDC tasks. This makes MolGPS the model with the most SOTA entries in the TDC leaderboard followed by MapLight+GNN [46] that established SOTA on 5 TDC tasks and 7 other methods that are SOTA for at least one TDC task. [^3] We report scores from the TDC leaderboards extracted on March 15, 2024. SOTA on TDC is established by a group of 8 different models, namely Chemprop-RDKit [78], MapLight, MapLight+GNN [46], BaseBoosting [27], CFA [49], SimGCN, ZairaChem [48] and ContextPred [85]. For MoleculeNet, we use the same split as GraphMVP [35], thus report the best results from their table. Similarly, compared to previous best methods on the Polaris and MoleculeNet benchmarks, we observe that our model is significantly better (often by large margins) for all but one downstream task. We primarily attribute the large-scale success to width scaling up to 3B parameters and the integration of phenomics data for pretraining. Figure 5 shows the normalized performance of MolGPS for the TDC benchmark, where it performs comparable to the best model per task for 1B parameters and clearly outperforms that baseline for 3B parameters. This is remarkable recalling again that this score is derived from the best scoring method per task of the benchmark collection, while we use a single method for all tasks. We finally note that further comparison of MolGPS to several foundation models that rely on self-supervised pretraining can be found in Appendix F.
@@ 5 Conclusion
In this paper, we studied the scalability of GNN models including message-passing networks, graph Transformers and hybrid architectures on the largest public collection of 2D molecules for the tasks of molecular property prediction. We showed major performance gains from the growing amount of parameters, data and compute, both on the original test set and on downstream finetuning. Importantly, our models benefit tremendously from the increasing scale of width, number of molecules, and number of labels. Our largest 3B parameter models, including MPNN++, Transformer, and GPS++, continue to scale favourably. More importantly, we demonstrate a consistent performance improvement on downstream property prediction tasks via finetuning and probing as we scale model and data size. Finally, we derive MolGPS, a powerful foundational model based on a multi-fingerprint probing approach that can be used to navigate the chemical space, establishing state-of-the-art on 26 out of 38 highly competitive downstream tasks. We hope that our work paves the way for the development of foundational GNNs and new architectures with applications in pharmaceutical advancements and drug discovery.
Future Work.
@@
While our study demonstrates the benefits of increasing number of parameters far greater than prior work, there are still orders of magnitude before we reach a general-purpose foundational model of molecules. Our analysis is restricted to the effect of number of parameters and molecules during pretraining and finetuning stages. Future work would aim to uncover additional aspects of GNN training such as the increasing complexity of aggregation functions and their effect on scaling properties. It will also be important to bridge current limitations for training large GNNs for molecules related to the expensive featurization as graphs, fast data loading and methods that unlock even larger models such as model parallelism.
Broader Impact.
@@
We foresee positive impacts of GNNs in areas of drug discovery, pharmaceutical advancements and tackling rare diseases by studying their molecular configurations. On the other hand, such models could also be used for harmful purposes such as developing chemical weapons and biohazards. We note that the usage of GNN models for such applications is less likely.
Appendix A Related Work
Foundation Models for Molecules. Recent work has relied on foundation models as a generalist class of models for sequential modelling [82, 33] as well as knowledge retrieval [17]. Within molecular drug discovery, recent works rely on structured models of ligands [40]. Luo et al. [37] and Moret et al. [41] study a general model for protein synthesis. Rao et al. [54] construct a self-attention driven architecture for contact prediction. Madani et al. [38] learn to generate a family of functional proteins. Nijkamp et al. [45] present a class of protein-pretrained language models. Similarly, Méndez-Lucio et al. [43] study binding interactions between different assays at the molecule-molecule interaction level. While many models focus on the design of molecules, a recent class of methods has also focused on properties of molecules [7]. Our study follows up by exploring similar molecular tasks for property prediction.
We elaborate more on additional advancements in foundational models making use of molecular graphs. Recent works have argued that the use of high-capacity models will be a significant boon to scientific discovery tasks [69]. Of specific interest are tasks in the quantum and molecular discovery paradigms [84] which demand domain-specific expertise such as knowledge of structure, provision of additional inductive biases and large data requirements. Towards this hypothesis, Fifty et al. [16] present an in-context learning framework for molecular property prediction without explicitly using a meta learning procedure. This leads to a general algorithm capable of discovering high-level structures from a pretraining sample set. Guo et al. [20] propose a similar framework making use of few-shot learning techniques resulting in a sample-efficient learning procedure. Xu et al. [75] present an alternative approach by modelling the full graphical structure of molecules across different property prediction tasks. Although effective, modelling the entire graph results in a computationally intensive learning procedure. Finally, Cao et al. [11] scale up learning to larger graph sizes by segmenting graph neighborhoods on the fly. An ad-hoc partitioning procedure is employed and interleaved with the learning phase in order to accelerate learning on larger and dense graphical clusters.
Architecture Design. Recent methods in graph architecture design focus on attending to structural information across nodes [44]. Of specific interest are graph Transformer networks which extract node as well as edge information by composing sequential attention modules over graph readouts [83]. In parallel, graph attention networks model attention weights across edges of a graph [67].
While attention mechanisms have demonstrated modern progress, traditional architectures such as neural message passing [18]. On one hand, message passing provides a rich and expressive framework for constructing representations. Godwin et al. [19] study regularization based on noisy nodes for the task of molecular property prediction. Provision of noise imputation in node-level features leads to simple and expressive method for tackling sparse molecular graphs. Graph bootstrapping [63] allows prior architectures to scale up to larger and complex graphs for representation learning. Our exploration of different architectures is aligned with the aforesaid works in literature, and with recent trends towards Transformers in related machine learning fields.
Scaling Laws. Recent work in model scaling has demonstrated that performant models follow a power law relationship between their parameter sizes and performance on new data samples [28]. Additionally, this relationship holds during the finetuning stage [25], thus indicating a strong reliance on model parameters. Bahri et al. [4] explain this power law fit by observing learning as moving on a smooth data manifold. Aghajanyan et al. [1] study the power law fit for mixed modality generative models, indicating that the scaling behavior is modality agnostic across various datasets. The result hints at the generality of scaling across different domains and applications, which can be extended to the study of scaling laws towards different training regimes (such as finetuning, downstream transfer and inference) and different problem domains (vision, language, audio, generative modelling, etc.) Our exploration of scaling behaviors in graph networks is motivated by the aforesaid directions.
Expressivity of GNNs. Prior work highlights that GNN architectures are limited in their expressivity to distinguish between graphs of similar node arrangements but different geometrical structures [72]. Various works indicate this as a consequence of aggregation functions and other design factors involved in GNN training [21]. On the other hand, recent work argues that only specific architectures are found robust to over-smoothing when building latent representations [14]. For instance, graph Transformers exhibit over-smoothing robustness as they utilize strong inductive biases such as attention. Xu et al. [73] connect the limited expressivity of GNNs with their ability to extrapolate on simpler tasks. Contrary to multi-layer networks, GNNs struggle to extrapolate on simpler tasks but show promise for improvement. Morris et al. [42] aim to tackle over-smoothing by building higher-order GNN architectures capable of capturing intricate node characteristics in their deeper layers. Finally, Ying et al. [81] present the differentiable pooling module capable of pooling neighboring node features which aid in reducing noise across layer representations.
Appendix B Experimental Details
B.1 Pretraining